A man walk 5kms due east and then 10kms due north. Find his
Total distance covered
Resultant displacement
1
Expert's answer
2019-08-22T09:25:53-0400

Total distance covered:


"d=5+10=15\ km"

Resultant displacement. Magnitude:


"s=\sqrt{5^2+10^2}=11.2\ km"

Direction:


"\theta=\arctan{\frac{5}{10}}=26.6\degree \ east \ of\ north"